OpenStudy (anonymous):

How many solutions does the equation 4x + 8 = –2 – 2 + 5x have? Two One None Infinitely many

OpenStudy (kidrah69):

combine like terms

OpenStudy (kidrah69):

-2-2=?

OpenStudy (anonymous):

i am just not quit sure if im doing it right so i am asking for help sorry if it is anoying

OpenStudy (kidrah69):

it okie :) thats why we are here

OpenStudy (kidrah69):

-2-2=?

OpenStudy (anonymous):

-4

OpenStudy (kidrah69):

Good now we have \[4x+8=-4+5x\] subtract 4x from both sides \[0x+8=-4+1x\]

OpenStudy (kidrah69):

add 4 again to both sides

OpenStudy (kidrah69):

8+4=?

OpenStudy (anonymous):

12

OpenStudy (kidrah69):

12/1=?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

4 + 8 = 1x remember 1x = 1

OpenStudy (anonymous):

i mean 1x=x

OpenStudy (kidrah69):

8+4=?

OpenStudy (kidrah69):

12

OpenStudy (kidrah69):

not 1

OpenStudy (anonymous):

x= 8+4

OpenStudy (anonymous):

so is it 12 or not

OpenStudy (kidrah69):

yeah so im right

OpenStudy (kidrah69):

12/1x=12

OpenStudy (kidrah69):

x=12

OpenStudy (anonymous):

so is the question infentley many?

OpenStudy (kidrah69):

well x equals a number

OpenStudy (kidrah69):

so it cant be infinate

OpenStudy (anonymous):

no since x = to only 12 then no

OpenStudy (anonymous):

so what will it be?

OpenStudy (kidrah69):

U tell me

OpenStudy (anonymous):

x = to only one number, not two numbers, and not infinite amount of numbers

OpenStudy (anonymous):

so there is only one

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

thanks!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!1

OpenStudy (anonymous):

np! anymore?

OpenStudy (anonymous):

yup!!